# I Introduction

### v1

@@ III.2 Sample robustness
We show the constant speed motion in the time histories is also consistently observed for three different values of = 6.35, 6.15 and 6.0 g/L in Fig. 4 (a). For these different samples, there is a good least squares fit of a straight line with the same accuracy as reported above for each time history which indicates that a constant velocity approximation is reasonable for the dynamics of the rising sphere towards a free surface. The terminal speed of the sphere can be estimated from the slope of a least squares fitted line through and an estimate of Fig. 4 (a) made in each case. The estimated value of can be used to rescale the time history plots and a convincing collapse of the data is found as shown over four orders of magnitude. Each measurement indicates an exponential dependence of on . This trend is shown in Fig. 4(b) plotted as a function of buoyancy stress . As expected, based on previous work [18], for each sample, the characteristic stress of the is slightly different.
We conclude that despite the constant speed motion of the sphere, the fluid medium is not Newtonian: there is an exponential drop in effective viscosity as the buoyancy is increased. Specifically, the log-linear plot of the slopes of the versus plots shown in Fig 4 indicates that the speed of the sphere depends exponentially on the buoyancy stress with stress factor , which varies between to Pa. This finding is consistent with previous results for stress dependence of creep at both local [17] and global [18] levels for both spheres and cylinders.
@@
Smaller diameter buoyant spheres are used to test the robustness of the results. These include a hollow plastic sphere of mm in diameter, which is filled with different amounts of water to reach the desired weight. Hollow PTFE, delrin and polypropylene spheres of mm, mm and mm are used. These intruders are used with hydrogel samples at and g/L. We find that the rising dynamics for the smaller diameters are in agreement with the data for the ping-pong balls over the entire range of buoyant stresses investigated as shown in Fig. 5. We conclude from these results that finite size effects of the spheres are not significant.
@@ IV Flow structure within the hydrogel suspension
The constant speed motion of the rising intruder begs the question if the flow field around the slowly moving sphere bears similarity to a Stokes flow. To investigate this, we perform experiments with ping pong balls rising to a free surface in a hydrogel suspension in the square container illuminated with a pair of green laser sheets as shown in Fig. 6. We analyze the measured flow fields around the sphere using the software PIVlab [30] applied to the videos recorded with the high-speed camera. The videos were pre-processed using intensity capping to reduce the noise from the brightest spots which arise from scattering of the laser illumination by impurities in the hydrogel suspension. We also remove the image of the ping pong ball from the region of interest for the PIV analysis using a mask.
@@
Finally, we remove the few outliers (which is a result of out-of-plane motion or poorly lit areas) through two velocity validation algorithms, automatically implemented in PIVlab. We first filter the velocity vectors based on the magnitude of each component, requiring that it does not deviate more than 8 standard deviations from the mean, and then we apply the local median filter [33]. This procedure yields a Valid Detection Probability (VDP) of higher than 98%. A typical PIV result from an experimental video is available as Supplementary Information.
The resulting PIV fields are averaged within a domain that co-moves with the steadily rising sphere obtained over 700 frames. The velocity field representing left-right motion is shown in Fig. 6 (b). The quadrupolar flow structure as required by mass conservation can be clearly distinguished, but there is a weak vertical asymmetry in the profile. The velocity component parallel to gravity is shown in Fig. 6 (c). The colour scheme highlights that the upward speed of the sphere is significantly higher than the return downward speed of the water in the hydrogels. This is reasonable since the downward counterflow required by conservation of mass is spread over a wide area around the sphere. To highlight the structure of the additional flow field induced by the shear between the sphere and the hydrogel sample, we constrain the colour scale in Fig. 6 (d). There is a large wake flow dragged upwards with the sphere, and there is also a large volume of hydrogel particles pushed ahead of the sphere. However, adjacent to the sphere, the hydrogels are mainly stationary, even though a no-slip boundary condition exists between the water and the sphere. Combining the velocity components into allows for a comparison of PIV results with tracking measurements and of the equatorial velocity component with the Stokes solution, Fig. 6 (e). We observe that the hydrogel fluid velocity profile has a very short decay length when compared with the long range Stokes solution.
Implicit in our arguments thus far is that the medium flows in a continuous way. In practice our sample comprises a mixture of polydisperse, deformable, hydrogel beads and water which could give rise to discrete displacement steps from e.g. stick-slip or other intermittent dynamics. We show such intermittency cannot be detected, even it high spatiotemporal resolution. In Fig. 7 the results of an experiment on the motion of a ping-pong ball recorded using the high speed imaging system described above that there is no evidence for discrete steps in the motion. Note that these results were obtained using the square container setup. The same qualitative behaviour was observed in separate experiments with a circular container, indicating that side-wall effects on the motion are also not significant.
@@ V Temperature and Diffusion effects
The constant speed behaviour suggests that a water-based lubrication mechanism may play a role in the dynamics of the rising sphere. We can test this hypothesis by measuring the effective viscosity of a sample under varying environmental temperature. The sample temperature is recorded in all experiments and the buoyancy stress is kept fixed for this particular investigation. We plot in Fig. 8 the effective viscosity of a hydrogel sample versus temperature with g/L with a buoyancy stress of Pa. The temperature range covered is to °C. Also included in Fig.8 are data points for water taken from tabulated values [34]. The data sets are normalised to be unity at °C. The fitted lines have an Arrhenius form; it can be seen that the hydrogel sample has a significantly stronger dependence on temperature than water. We therefore conclude that a simple water-based lubrication mechanism, for example in the observed slip between the sphere and the hydrogel sample found here and reported elsewhere [29], cannot be the only factor that plays a role in the effective viscosity of the two-phase medium. We can speculate on other temperature dependent mechanisms that contribute to the effective viscosity of the hydrogel sample. One of these can be the temperature dependence of the stiffness of the hydrogels [35], which controls their poroelastic behavior [36]. However, poroelastic effects typically have a timescale that is the ratio of the stiffness and the viscosity, which for the system observed here both decrease with temperature, partially compensating this effect.
@@
Another local mechanism that can drive motion of an intruder is local particle diffusion in the hydrogel suspension. It is thus interesting to understand how fast the intruder moves with respect to diffusive processes in the hydrogel suspension. The intruder motion is of the order of to m/s. There are two types of diffusive processes involved. At the molecular level motion within the water is one, but also the hydrogel particles close to the sphere are affected by passage of the intruder [37]. More intensive Particle Tracking Velocimetry could aid in extracting the mean squared displacment of the hydrogel particles. With the intruder speed and a characteristic lengthscale , one can estimate a Peclet number . From the available dyed tracer particle measurement we estimate that hydrogel particles do not diffuse more than a millimeter during the course of an experiment, putting an upper bound on their mean squared displacement m2/s. The diffusion constant of water is around m2/s, suggesting that the Peclet number should be larger than one for both the linear and sublinear scenarios. This simplified interpretation may be more subtle in practice: for example we expect to be dependent, as the diffusive motion of the hydrogels is driven not by thermal motion of the solvent molecules, but by the movement of the intruder [37], involving a self-fluidization process that is commonly observed to be present for dry granular materials [38, 39, 40]. The typical lengthscale involved is also not easily identified. We thus see several avenues for future research in exploring the role of microscopic mechanisms that may contribute to drive the slow flow in hydrogel suspensions.
@@ VI End boundary conditions
As outlined above in the discussion of the results in Fig. 1, the qualitative difference in the time series plots for cases with and without a lid is clear. At the same value of and buoyancy the sphere rises with a linear distance-time relationship with a free surface and is sub-linear, tending towards , when the lid is present. Note that a dependence of distance travelled on is qualitatively different to the end effects in falling-bob experiments on viscous Newtonian flows [6].
Thus, we systematically explore the effects of a lid on behavior. The results of tests with ping pong balls of various weights are shown in Fig. 10 (a). As shown in the example data from Fig. 9, we observe a clear sublinear time-dependent behavior of the displacement of the sphere. To extract the time dependence of the displacement of the rising sphere, we again fit to each data set
Here, is again a prefactor that sets the speed of motion. Despite the non-Newtonian nature of the flow behavior, we refer to it as the effective viscosity. We can verify which rescaling coefficient provides the best collapse of the data; plotting results in a linear time dependence. We find that provides the best collapse of the data and show this result in Fig. 10b. We verify the exponential dependence of prefactor that we use to characterize the strength of the sublinear behavior in Fig. 10c. Since the same hydrogel sample was used here, we can compare the exponential dependence on the buoyancy stress for the sublinear case to the case of the rising sphere towards the free surface as discussed in Sec. III.1 (dash-dotted line). The data for the sublinear case clearly follows the same scaling of , with the same Pa as used for the free surface case. We conclude that motion of a sphere towards a rigid boundary is sublinear by way of contrast with the motion towards a free surface, which is linear.
@@ VI.2 Sphere moving towards a rigid base — comparison with previous work
In previous work [17, 18] we performed experiments on sinking intruders that moved towards the rigid base of the container. These additional data enable us to compare the observed time dependence with the cases discussed in this manuscript. We extract the data from the open repositories, e.g. [41] and perform the same fitting procedure to estimate the behavior for these data sets.
We can extract the exponent for which a power law model provides the best fit. We note that there is no canonical definition of a “best” fit, as fit quality as measured by a reduced depends also on weights associated to data, and even with an optimal , systematic deviations in the residual of the fit are neglected. Our choice for “best” is hence simply the best goodness of fit estimate as returned by the MATLAB-based fitting routine.
Applying this same standard across all data sets from two previous publications and the current one, yields the results as shown in Fig. 11, with over 300 sets fitted. The statistics of the distribtion of for each case of “rigid” boundary, “floating lid” and “free” surface is indicated with a violin plot, including a separate scatter plot for the mean and standard deviation of the distribution (error bar). We observe a clear clustering of exponents around 0.5 for all the previously published data, in which an intruder was sinking towards a rigid base. In the current study, some data was obtained with an intruder rising towards a rigid lid floating on the hydrogel packing; the exponents for these experiments center around 0.7. The exponents for the creeping flow measurements for intruders that move towards a free boundary clearly center around , confirming again the strong influence of the boundary on the flow behavior of the hydrogel suspension. Lastly, for the data sets of the previous works we can obtain a satisfactory collapse as obtained for the current data as shown in Fig. 10 (b). The temporal resolution in our previous work was lower than used here, but the exponent which provides a satisfactory collapse for these datasets is for data from [17, 18] albeit with increased scatter. All data and analysis scripts for all our work can be found in open data repositories [41].
@@ VI.3 Rising towards a surface with a mesh
We show in Fig. 12 four time histories with hydrogel density g/L. Two of the time histories were measured with a free surface and two with a metal mesh which sits mm below the surface of the packing. The mesh prevents the hydrogel particles at the surface from moving during the transit of the ping-pong ball, yet it is sufficiently porous that water will flow relatively unimpeded. This boundary condition can is rough and hence suppresses vertical and radial motion of the hydrogels at the top of the packing. However, the mesh is also very light. The results indicate that the mesh has very little effect on the constant speed motion of the ping-pong ball. Moreover, the surface of the sample remains flat during the motion of the ball, suggesting that the constant speed motion is not an end effect imposed by having a surface structure imposed, suggesting that flexible rough lids would also not create sublinear motion. We cannot exclude the possibility that a smooth yet flexible boundary such as the surface tension boundary conditions from [17] can nevertheless be essential to create sublinear motion.
@@
Experiments were performed to investigate whether the constant speed behavior exists for heavy spheres that sink towards a “free” boundary formed under the hydrogel suspension by injecting a heavy immiscible fluid at the bottom of the container. In these experiments, the hydrogel suspension or was created on top of a cm liquid layer of Galden 135 fluid. The hydrogel suspension was made with deionized water, in which the hydrogel particles swell more easily. This accounts for the lower value of used here in comparison to the typical values used in the ping-pong ball experiments. We confirmed that this lower value of produced a sample in which the entire water volume is densely packed with hydrogel spheres.
A water-filled ping pong ball is used in the initial experiments. Hence it is negatively buoyant in the hydrogel suspension, but positively buoyant in the heavy Galden liquid. The motion of the filled ping-pong ball achieves a constant velocity while fully immersed in the hydrogel suspension but slows to a standstill at the interface of the hydrogel suspension and the Galden. The observations are found to be repeatable. The second set of experiments involves a glass sphere (kgm-3), which is negatively buoyant in both the hydrogel suspension and Galden. The hydrogel suspension used in the second experiment is slightly denser to set the sinking time to be consistent with previous observations. Despite the differences between the two sinking experiments, the spheres reached constant velocity as indicated in Fig. 13 (a). Rescaling the displacement data with the effective viscosity of the respective suspension-intruder pairs collapses the data in a straight line with slope , confirming once more that even if the motion is in the direction of gravity, an intruder driven by a constant stress moving towards a free boundary of a hydrogel suspension maintains a constant speed.

### v2

@@ III.2 Sample robustness
We show the constant speed motion in the time histories is also consistently observed for three different values of and g/L in Fig. 4 (a). For these samples, there are good least squares fits of a straight lines with the same accuracy as reported above. This indicates a constant velocity approximation is an appropriate descriptor for the dynamics of the sphere rising towards a free surface. The terminal speed of the sphere can be estimated from the slope of the least squares fitted line through and an estimate of Fig. 4 (a) is made in each case. The estimated value of can be used to rescale the time history plots and a convincing collapse of the data is found as shown over four orders of magnitude. Each measurement indicates an exponential dependence of on . This trend is shown in Fig. 4(b) plotted as a function of buoyancy stress . As expected, based on previous work [22], for each sample, the characteristic stress of the is slightly different.
We conclude that despite the constant speed motion of the sphere, the fluid medium is not Newtonian: there is an exponential drop in effective viscosity as the buoyancy is increased. Specifically, the log-linear plot of the slopes of the versus plots shown in Fig 4 indicates that the propagation speed of the sphere depends exponentially on the buoyancy stress with stress factor , which varies between to Pa. This finding is consistent with previous results for stress dependence of creep at both local [21] and global [22] levels for both spheres and cylinders.
@@
Smaller diameter buoyant spheres are used to test the robustness of the results. These include a hollow plastic sphere of mm in diameter, which is filled with different amounts of water to reach the desired weight. Hollow PTFE, delrin and polypropylene spheres of mm, mm and mm are also employed. These intruders are used with hydrogel samples at and g/L. We find that the rising dynamics for the smaller diameters are in agreement with the data for the ping-pong balls over the entire range of buoyant stresses investigated as shown in Fig. 5. We conclude from these results that finite size effects of the spheres are not significant.
@@ IV Flow structure within the hydrogel suspension
The constant speed motion of the rising intruder begs the question if the flow field around the slowly moving sphere bears similarity to a Stokes flow. To investigate this, we perform experiments with ping pong balls rising towards a free surface in a hydrogel suspension in a square container illuminated by a pair of green laser sheets as shown schematically in Fig. 6. We analyze the measured flow fields around the sphere using the software PIVlab [36] applied to the videos recorded with the high-speed camera. The videos were pre-processed using intensity capping to reduce the noise from the brightest spots which arise from scattering of the laser illumination by impurities in the hydrogel suspension. We also remove the image of the ping pong ball from the region of interest for the PIV analysis using a mask.
@@
Finally, we remove the few outliers (which is a result of out-of-plane motion or poorly lit areas) through two velocity validation algorithms, automatically implemented in PIVlab. We first filter the velocity vectors based on the magnitude of each component, requiring that it does not deviate by more than 8 standard deviations from the mean, and then we apply the local median filter [39]. This procedure yields a Valid Detection Probability (VDP) of higher than 98%. A typical PIV result from an experimental video is available as Supplementary Information.
The resulting PIV fields are averaged within a domain that co-moves with the steadily rising sphere obtained over 700 frames. The velocity field representing left-right motion is shown in Fig. 6 (b). The quadrupolar flow structure as required by mass conservation can be clearly distinguished, and there is a weak vertical asymmetry in the profile. The velocity component parallel to gravity is shown in Fig. 6 (c). The colour scheme highlights that the upward speed of the sphere is significantly higher than the return downward speed of the water in the hydrogels. This is reasonable since the downward counterflow required by conservation of mass is spread over a wide area surrounding the sphere. To highlight the structure of the additional flow field induced by the shear between the sphere and the hydrogel sample, we constrain the colour scale in Fig. 6 (d). There is a significant wake flow dragged upwards with the sphere, and there is also a large volume of hydrogel particles pushed ahead of the sphere. However, adjacent to the sphere, the hydrogels are mainly stationary, even though a no-slip boundary condition exists between the water and the sphere. Our measurements are accurate up to a distance of mm (approximately the size of a particle) from the spherical intruder. Combining the velocity components into permits a comparison of PIV results with tracking measurements and the equatorial velocity component with the Stokes solution, Fig. 6 (e). We observe that the hydrogel fluid velocity profile has a very short decay length when compared with the long range Stokes solution.
Implicit in our arguments thus far is that the medium flows in a continuous way. In practice our sample comprises a mixture of polydisperse, deformable, porous, hydrogels and water which could give rise to discrete displacement steps from e.g. stick-slip or other intermittent dynamics. We are unable to detect intermittency when we use higher spatiotemporal resolution. In Fig. 7 the results of an experiment on the motion of a ping-pong ball recorded using the high speed imaging system described above show there is no evidence for discrete steps in the motion. Note that these results were obtained using the square container setup. The same qualitative behaviour was observed in separate experiments with a circular container, indicating that side-wall effects on the motion are also not significant.
@@ V Temperature and Diffusion effects
The constant speed behaviour suggests that a water-based lubrication mechanism may play a role in the dynamics of the rising sphere. We can test this hypothesis by measuring the effective viscosity of a sample under varying environmental temperature. The sample temperature is recorded in all experiments and the buoyancy stress is kept fixed for this particular investigation. We plot in Fig. 8 the effective viscosity of a hydrogel sample versus temperature with g/L with a buoyancy stress of Pa. The temperature range covered is to °C. Also included in Fig.8 are data points for water taken from tabulated values [40]. The data sets are normalised to be unity at °C. The fitted lines have an Arrhenius form; it can be seen that the hydrogel sample has a significantly stronger dependence on temperature than water. We conclude that a simple water-based lubrication mechanism, for example in the observed slip between the sphere and the hydrogel sample found here and reported elsewhere [35], cannot be the only factor that plays a role in the effective viscosity of the two-phase medium. We speculate that other temperature dependent mechanisms contribute to the effective viscosity of the hydrogel sample. One of these could be the temperature dependence of the stiffness of the hydrogels [41], which controls their poroelastic behavior [42]. However, poroelastic effects typically have a timescale that is the ratio of the stiffness and the viscosity, which for the system observed here both decrease with temperature, partially compensating this effect.
@@
Another local mechanism that can drive motion of an intruder is local particle diffusion in the hydrogel suspension. It is thus interesting to understand how fast the intruder moves with respect to diffusive processes in the hydrogel suspension. The intruder motion is of the order of to m/s. There are two types of diffusive processes involved. At the molecular level motion within the water is one, but also the hydrogel particles close to the sphere are affected by passage of the intruder [43]. More intensive Particle Tracking Velocimetry could aid in extracting the mean squared displacment of the hydrogel particles. With the intruder speed and a characteristic lengthscale , one can estimate a Peclet number . From the available dyed tracer particle measurement we estimate that hydrogel particles do not diffuse more than a millimeter during the course of an experiment, putting an upper bound on their mean squared displacement m2/s. The diffusion constant of water is around m2/s, suggesting that the Peclet number should be larger than one for both the linear and sublinear scenarios. This simplified interpretation may be more subtle in practice: for example we expect to be dependent, as the diffusive motion of the hydrogels is driven not by thermal motion of the solvent molecules, but by the movement of the intruder [43], involving a self-fluidization process that is commonly observed to be present for dry granular materials [44, 45, 46]. The typical lengthscale involved is also not easily identified. We thus see several avenues for future research to explore the role of microscopic mechanisms that may contribute to driving the slow flow in hydrogel suspensions.
@@ VI End boundary conditions
As outlined above in the discussion of the results in Fig. 1, the qualitative difference in the time series plots for cases with and without a lid is clear. At the same value of and buoyancy, the sphere rises with a linear distance-time relationship with a free surface . However, it is sub-linear, approximating previously observed scaling, when the lid is present. Note that a dependence of distance travelled on is qualitatively different to the end effects in falling-bob experiments on viscous Newtonian flows [7].
The sublinear dynamics observed suggests a systematic exploratiom of the effects of a lid on behavior is required. The results of tests with ping pong balls of various buoyancies are shown in Fig. 10 (a). As shown in the example data from Fig. 9, we observe a clear sublinear time-dependent behavior of the displacement of the sphere. In order to extract the time dependence of the displacement of the rising sphere, we again fit to each data set
Here, is a prefactor set by the speed of motion. Despite the non-Newtonian nature of the flow behavior and the unusual dimensions of , we refer to it as the effective viscosity. We verify which rescaling coefficient provides the best collapse of the data; plotting results in a linear time dependence. We find that provides the best collapse of the data and show this result in Fig. 10b. We verify the exponential dependence of prefactor that we use to characterise the strength of the sublinear behavior in Fig. 10c. Since the same hydrogel sample was used here, we can compare the exponential dependence on the buoyancy stress for the sublinear case with the case of the sphere rising towards a free surface as discussed in Sec. III.1 (dash-dotted line). The data for the sublinear case clearly follows the same scaling of , with the same Pa as used for the free surface case. We cannot expect a perfect match in the functional form, as absorbs some of the physical prefactors that we could separate out in the constant velocity case. We conclude that motion of a sphere towards a rigid boundary is sublinear by way of contrast with the motion towards a free surface, which is linear.
@@ VI.2 Sphere moving towards a rigid base — comparison with previous work
In previous work [21, 22] we performed experiments on sinking intruders that moved towards the rigid base of the container. These additional data enable us to compare the observed time dependence with the cases discussed in this manuscript. We extract the data from the open repositories, e.g. [47] and perform the same fitting procedure to estimate the behavior for these data sets. To allow for a systematic comparison of our results, we publish our kinetic data and analysis scripts in an open repository [48].
We can extract the exponent for which a power law model provides the best fit. We note that there is no canonical definition of a “best” fit, as fit quality as measured by a reduced depends also on weights associated with data, and even with an optimal , systematic deviations in the residual of the fit are neglected. Our choice for “best” is hence simply the best goodness of fit estimate as returned by the MATLAB-based fitting routine.
Applying this same standard across all data sets from two previous publications and the current one, yields the results as shown in Fig. 11, with over 300 sets fitted. The statistics of the distribtion of for each case of “rigid” boundary, “floating lid” and “free” surface is indicated with a violin plot, including a separate scatter plot for the mean and standard deviation of the distribution (error bar). We observe a clear clustering of exponents around 0.5 for all the previously published data, in which an intruder was sinking towards a rigid base. In the current study, some data was obtained with an intruder rising towards a rigid lid floating on the hydrogel packing; the exponents for these experiments center around 0.7. The exponents for the creeping flow measurements for intruders that move towards a free boundary clearly center around , confirming again the strong influence of the boundary on the flow behavior of the hydrogel suspension. Lastly, for the data sets of the previous investigations we can obtain a satisfactory collapse as obtained for the current data as shown in Fig. 10 (b). The temporal resolution in our previous work was lower than used here, but the exponent which provides a satisfactory collapse for these datasets is for data from [21, 22] albeit with increased scatter. All data and analysis scripts for all our work can be found in open data repositories [47].
@@ VI.3 Rising towards a surface with a mesh
We show in Fig. 12 four time histories with hydrogel density g/L. Two of the time histories were measured with a free surface and two with a metal mesh which sits mm below the surface of the packing. The mesh prevents the hydrogel particles at the surface from moving during the transit of the ping-pong ball, yet it is sufficiently porous that water will flow relatively unimpeded. This boundary condition is rough and hence suppresses vertical and radial motion of the hydrogels at the top of the packing. However, the mesh is also very light. The results indicate that the mesh has very little effect on the constant speed motion of the ping-pong ball. Moreover, the surface of the sample remains flat during the motion of the ball, indicating that the constant speed motion is not an end effect imposed by having a surface structure imposed, suggesting that flexible rough lids would also not create sublinear motion. We cannot exclude the possibility that a smooth yet flexible boundary such as the surface tension boundary conditions from [21] can nevertheless be essential to create sublinear motion.
@@
Experiments were performed to investigate whether the constant speed behavior exists for heavy spheres that sink towards a “free” boundary formed under the hydrogel suspension by placing a heavy immiscible fluid at the bottom of the container. In these experiments, the hydrogel suspension or was created on top of a cm liquid layer of Galden 135 fluid. The hydrogel suspension was made with deionised water, in which the hydrogel particles swell more easily. This accounts for the lower value of used here in comparison to the typical values used in the ping-pong ball experiments. We confirmed that this lower value of produced a sample in which the entire water volume is densely packed with hydrogel spheres.
A water-filled ping pong ball is used in the initial experiments. Hence, it is negatively buoyant in the hydrogel suspension, but positively buoyant in the heavy Galden liquid. The motion of the filled ping-pong ball achieves a constant velocity while fully immersed in the hydrogel suspension but slows to a standstill at the interface of the hydrogel suspension and the Galden. The observations are found to be repeatable. The second set of experiments involves a glass sphere (kgm-3), which is negatively buoyant in both the hydrogel suspension and Galden. The hydrogel suspension used in the second experiment is slightly denser to set the sinking time to be consistent with previous observations. Despite the differences between the two sinking experiments, the sphere reaches constant velocity as indicated in Fig. 13 (a). Rescaling the displacement data with the effective viscosity of the respective suspension-intruder pairs collapses the data in a straight line with slope , confirming once more that even if the motion is in the direction of gravity, an intruder driven by a constant stress moving towards a free boundary of a hydrogel suspension maintains a constant speed.